# Exponentially Consistent Nonparametric Linkage-Based Clustering of Data Sequences

### v2

@@ I Introduction
$N$ Stopping time of clustering algorithm
@@ IV Proposed Sequential nonparametric clustering algorithm
$M_{b}(i,j,n)=$$\left\{\left(\frac{n-1}{n}\right)^{2}M_{b}^{2}(i,j,n-1)\right.$
$+\frac{1}{n^{2}}\sum_{l=1}^{n}h\left(X_{l}^{(i)},X_{n}^{(i)},X_{l}^{(j)},X_{n}^{(j)}\right)$
$\left.+\frac{1}{n^{2}}\sum_{m=1}^{n-1}h\left(X_{n}^{(i)},X_{l}^{(i)},X_{n}^{(j)},X_{l}^{(j)}\right)\right\}^{\frac{1}{2}}$
@@
[…]
5: while $\textbf{while }{\Gamma_{n}<\frac{C}{\sqrt{n}}}$
@@ IV-A Analysis of SLINK-SEQ
=[$=\mathbb{P}\left[\{N>n\}\cap E_{n}\right]+\mathbb{P}\left[\{N>n\}\cap E_{n}^{c}\right]$
@@
[…]
@@
[$\mathbb{P}\left[\{N>n\}\cap E_{n}^{c}\right]\leq M^{2}\left(2\exp\left(\frac{-n(\delta d_{H})^{2}}{16\mathbb{G}}\right)\right),$
@@
[…]
@@
[…]
@@ IV-B Complexity of the SLINK-SEQ
  • Update the pairwise distances between the data sequences. There are $M(M-1)/2$ distances to be updated. The MMD can be updated using the sequential update in $O(n)$ computations. Therefore, the overall complexity of this step is of the order $O(nM^{2})$.
  • Updating the clusters using SLINK: SLINK starts from $M$ clusters and reduces the number of clusters one at a time to $K$ clusters. Therefore, there are $M-K$ steps. In each step of SLINK, the distance of the merged cluster with the other clusters needs to be updated. This would require $O(M-i)$ computations at step $i$. The overall complexity of this step is $O(M^{2})$.
Overall, the complexity grows at most quadratically in the number of data sequences $M$. Further simplification could be studied in future work.
@@ V Simulation Results
We now study the performance of SLINK and SLINK-SEQ using simulations. For the FSS setting, we illustrate that SLINK is exponentially consistent under the condition $d_{I}<d_{H}$. We also compare SLINK with $k$-medoids clustering (KMed) in [19]. Then, we compare SLINK with SLINK-SEQ and show how sequential clustering can provide improved performance. In the simulations, we run the clustering algorithms for multiple realizations of the data sequences and compare the final clustering output to the true clusters. An error is declared if the clustering output does not exactly match the true clusrers. We estimate $P_{e}$ by taking the ratio of the number of errors to the number of realizations simulated. In the case of sequential algorithms, we also determine the number of samples at which the algorithm makes a decision on the clusters for each realization. $\mathbb{E}[N]$ is estimated as the average number of samples taken over all the realizations.
@@ V-A Simulation settings
First, we consider three examples to illustrate the difference between the distance assumptions between the clusters. Examples 1 and 2 have $d_{L}>d_{H}$ but $d_{I}<d_{H}$. Example 3 satisfies $d_{L}<d_{H}$ and has been used in [19]. In each example, the i.i.d. data sequences are generated according Gaussian distributions with unit variance and different means. The three examples are: (1) 2 clusters, means of the sequences in cluster 1 and cluster 2 are $\{0.4,0.55,0.7,0.85,1,1.15,1.3,1.45,1.6\}$ and $\{1.85,2,2.15\}$, (2) 2 clusters, means of the sequences in cluster 1 and cluster 2 are $\{0.7,0.85,1,1.15,1.3\}$ and $\{1.7,1.85,2,2.15,2.3\}$, and (3) 5 clusters, 5 sequences per cluster, mean of each sequence in cluster $k$ is $k-1$ for $k=1,2,3,4,5$.
For the above examples, $d_{I}$ is equal to the maximum gap between two neighbouring distributions within a cluster, i.e, for each distribution we need only the left and right neighbours to be closer than the distributions in the other clusters. Therefore, we can have big clusters that are separated by gaps that are small (as in Example 1 and 2). In fact, for these examples, clustering is achieved by identifying the largest gap between means. Table II shows the MMD and KSD distances $d_{L}$, $d_{H}$, and $d_{I}$ for each of the three examples. The distances are estimated using 10000 samples from each distribution. Clearly, we have $d_{I}<d_{H}$ in all three examples, while $d_{L}<d_{H}$ is true only in example 3.
@@
Next, we consider two more examples with non-Gaussian data sequences. There are 2 clusters with 3 data sequences each. Here, each data sequence is generated in an i.i.d. manner according to a Gaussian Mixture Model (GMM). In particular, each sequence is a mixture of 2 Gaussian distributions: $N(m_{1},1)$ with probability $0.7$, and $N(m_{2},1)$ with probability 0.3. For Example 4, the $(m_{1},m_{2})$ values for the sequences in the two clusters are as follows: {$(-0.5,0),(0,0.5),(0.5,1)$}, and {$(1.2,1.7),(1.7,2.2),(2.2,2.7)$}. For Example 5, the $(m_{1},m_{2})$ values for the sequences in the two clusters are as follows: {$(-0.5,0),(0,0.5),(0.5,1)$}, and {$(1.35,1.85),(1.85,2.35),(2.35,2.85)$}. It can be seen that the clusters are closer in Example 4 compared to Example 5. The distances $d_{L}$, $d_{I}$, and $d_{H}$ for these two examples are given in Table III. In both cases, $d_{L}>d_{H}$, but $d_{I}<d_{H}$. In particular, $d_{I}$ is only slightly lesser than $d_{H}$ in Example 4.
Finally, we consider an example using Modified National Institute of Standards and Technology (MNIST) data [29]. This data set contains about 6000 images each of handwritten numerals from 0 to 9 of size $28\times 28$. Using this data set, we create five data sequences of 1000 samples for each numeral from 0 to 9. Thus, we have 50 data sequences forming 10 clusters. Here, each sample in a data sequence has $28\times 28=784$ dimensions For this multi-dimensional sample case, we use the kernel function:
$k({\mathbf{x}},{\mathbf{y}})=e^{-||{\mathbf{x}}-{\mathbf{y}}||^{2}/2}.$
We also created another lower dimensional data set that reduces the 784 dimensions to 10 dimensions using principal component analysis.
@@ V-B FSS performance
Fig. 5 shows $\text{ln}(P_{e})$ versus the sample size $n$ for Example 1. The performance of SLINK is shown for both MMD (with kernel $k(x,y)=e^{-(x-y)^{2}/2}$) and KSD as the distance measures. It can be observed that SLINK is exponentially consistent, i.e., the $\text{ln}(P_{e})$ vs. $n$ plot is linearly decreasing, as expected. For this example, KMed is unable to find the true clusters in the FSS setting for any $n$. Although only the KSD case is shown for KMed, KMed with MMD also does not work. This is mainly due to the fact that one cluster is big while the other is small, and some distributions in the big cluster are closer to the medoid of the small cluster. However, SLINK can work even in this setting as $d_{I}<d_{H}$ is satisfied as long as the left and right neighbours for each distribution (in terms of the means of the Gaussian distributions) within the same cluster are closer than the distributions from the other clusters.
Figs. 6 and 7 show $\text{ln}(P_{e})$ versus the sample size $n$ for Examples 2 and 3. Both SLINK and KMed are exponentially consistent in the FSS setting for these examples. Even though $d_{L}>d_{H}$, KMed works for Example 2 since the clusters are of equal size and, therefore, all the distributions in each cluster are closer to the medoid of the correct cluster. For these two examples, KMed is able to perform better than SLINK, i.e., it achieves the same error probability at lower $n$. In particular, for example 2, KMed with MMD estimates can achieve $\text{ln}(P_{e})=-1.5$ with 100 samples, while SLINK with MMD estimates requires around 200 samples. However, SLINK can work for all problems with $d_{I}<d_{H}$, while KMed cannot. Fig. 7 also shows the performance of CLINK using both KSD and MMD estimates. It can be observed that for this example, CLINK is also exponentially consistent. It performs better than SLINK for lower $n$. However, CLINK does not work for Examples 1 and 2 since $d_{L}>d_{H}$ in these two examples.
@@ V-C Sequential clustering performance
Fig. 10 shows $\text{ln}(P_{e})$ versus $E[N]$ for SLINK-SEQ and SLINK FSS using MMD for the GMM Examples 4 and 5. Again, as in the previous examples, we observe that the proposed sequential clustering algorithm requires fewer expected number of samples than the FSS test for the same $P_{e}$. Finally, in Fig. 11, we show the performance of SLINK FSS and SLINK-SEQ for the MNIST data set. Again, the algorithms are exponentially consistent. The proposed SLINK-SEQ algorithm is able to successfully cluster the data sequences corresponding to the same number. We also observe that the performance with the 784-dimensional data is better than the performance with the 10-dimensional data obtained using principal component analysis.
@@
In Figs. 12 and 13, we show the performance of SLINK-SEQ for different threshold choices for Examples 1, 2 and 3, respectively. We use the threshold $C/n^{\alpha}$ with $\alpha=1/3,1/2,1$. We observe that for all cases, $\alpha=1/2$ results in better performance.
@@ Appendix A Concentration of MMD estimates
$P\left(\mathbb{MMD}\left(\mathrm{x}_{1},\mathrm{x}_{2}\right)\leq d_{0}\right)\leq 2\exp(\frac{-n(d_{H}-d_{0})^{2}}{16\mathbb{G}})$ (19)
@@
$P\left(\mathbb{MMD}\left(\mathrm{x}_{1},\mathrm{x}_{2}\right)\geq d_{0}\right)\leq 2\exp(\frac{-n(d_{0}-d_{I})^{2}}{16\mathbb{G}})$ (21)

### v4

@@ I Introduction
Overall, we demonstrate that, in the FSS setting, SLINK clustering of data sequences can achieve exponential consistency under a less restrictive condition than previously known, and subsequently, propose an exponentially consistent clustering algorithm, SLINK-SEQ, in the sequential setting.
@@
NN Stopping time of clustering algorithm
@@ IV Proposed Sequential nonparametric clustering algorithm
$M_{b}(i,j,n)=$ $\left\{\left(\frac{n-1}{n}\right)^{2}M_{b}^{2}(i,j,n-1)\right$. (6)
$+\frac{1}{n^{2}}\sum_{l=1}^{n}h\left(X_{l}^{(i)},X_{n}^{(i)},X_{l}^{(j)},X_{n}^{(j)}\right)\hskip-57.81621pt$
$\left.+\frac{1}{n^{2}}\sum_{m=1}^{n-1}h\left(X_{n}^{(i)},X_{l}^{(i)},X_{n}^{(j)},X_{l}^{(j)}\right)\right\}^{\frac{1}{2}}\hskip-57.81621pt$
@@
3: $\textbf{Initialize: }n\leftarrow 2$, 22 samples per sequence, $T_{2}=0$
4: $\textbf{Calculate distances: }\hat{d}(i,j,n)$ for $i<j$, $i,j\in\{1,2,\ldots,M\}$
5: $\textbf{while }{\Gamma_{n}<\frac{C}{\sqrt{n}}}$
@@ IV-A Analysis of SLINK-SEQ
$=\mathbb{P}\left[\{N>n\}\cap E_{n}\right]+\mathbb{P}\left[\{N>n\}\cap E_{n}^{c}\right]$
@@
$\text{(or)}~4\sqrt{\frac{\mathbb{G}}{n}}<\frac{d_{H}-T_{n}}{2}-\frac{\delta}{2}d_{H}$ (8)
$\text{(or)}~8\sqrt{\frac{\mathbb{G}}{n}}+\frac{C}{\sqrt{n}}<(1-\delta)d_{H}$
$\text{(or)}~n>\left(\frac{C+8\sqrt{\mathbb{G}}}{(1-\delta)d_{H}}\right)^{2}.$
@@
$\mathbb{P}\left[\{N>n\}\cap E_{n}^{c}\right]\leq M^{2}\left(2\exp\left(\frac{-n(\delta d_{H})^{2}}{16\mathbb{G}}\right)\right),$
@@
$\text{(or)}~4\sqrt{\frac{\mathbb{G}}{n}}<\frac{1-\delta}{2}T_{n}-\frac{d_{I}}{2}$
$\text{(or)}~8\sqrt{\frac{\mathbb{G}}{n}}<(1-\delta)\frac{C}{\sqrt{n}}-d_{I}$
$\text{(or)}~C>\frac{d_{I}\sqrt{n}+8\sqrt{\mathbb{G}}}{(1-\delta)}$
@@
$\text{(or)}~(1-\delta)C>d_{I}\left(\sqrt{\tilde{n}}+\frac{C}{(1-\delta)d_{H}}\right)+8\sqrt{\mathbb{G}}$
$\text{(or)}~C>\frac{d_{I}\sqrt{\tilde{n}}+8\sqrt{\mathbb{G}}}{(1-\delta)-\frac{d_{I}}{(1-\delta)d_{H}}}$ (16)
@@ IV-B Complexity of the SLINK-SEQ
  • Updating the pairwise distances between the data sequences: There are M(M1)/2M(M-1)/2 distances to be updated. The MMD can be updated using the sequential update in $O(n)$ kernel computations. The kernel can computed in $O(n_{\text{dim}})$ computations, where $n_{\text{dim}}$ is the dimension of each sample of a data sequence. Therefore, the overall complexity of this step is of the order $O(dn_{\text{dim}}M^{2})$.
  • Updating the clusters using SLINK: SLINK starts from $M$ clusters and reduces the number of clusters one at a time to $K$ clusters. Therefore, there are MKM-K steps. In each step of SLINK, the distance of the merged cluster with the other clusters needs to be updated. This would require O(Mi)O(M-i) computations at step $i$. The overall complexity of this step is $O(M^{2})$.
Overall, the complexity grows at most quadratically in the number of data sequences $M$ and linearly in the number of dimensions $n_{\text{dim}}$ of each sample in a data sequence. Further simplification could be studied in future work.
The cost of sequential MMD update in (6) grows as more data is observed, as it requires $O(n)$ complexity. Some simplifications of this by using only a limited window of prior samples in the MMD update could be considered to reduce this complexity. However, further investigation is required to identify the right window size. This issue and other related issues on the scalability of clustering algorithms have been studied in [29].
@@ V Simulation Results
We now study the performance of SLINK and SLINK-SEQ using simulations. [^8] The code used to generate the results in this paper are available at https://github.com/bhupenderee22s006/TSP_codes.git For the FSS setting, we illustrate that SLINK is exponentially consistent under the condition $d_{I}<d_{H}$. We also compare SLINK with $k$-medoids clustering (KMed) in [19]. Then, we compare SLINK with SLINK-SEQ and show how sequential clustering can provide improved performance. In the simulations, we run the clustering algorithms for multiple realizations of the data sequences and compare the final clustering output to the true clusters. An error is declared if the clustering output does not exactly match the true clusrers. We estimate $P_{e}$ by taking the ratio of the number of errors to the number of realizations simulated. In the case of sequential algorithms, we also determine the number of samples at which the algorithm makes a decision on the clusters for each realization. $\mathbb{E}[N]$ is estimated as the average number of samples taken over all the realizations.
@@ V-A Simulation settings
First, we consider three examples to illustrate the difference between the distance assumptions between the clusters. Examples 1 and 2 have $d_{L}>d_{H}$ but $d_{I}<d_{H}$. Example 3 satisfies $d_{L}<d_{H}$ and has been used in [19]. In each example, the i.i.d. data sequences are generated according Gaussian distributions with unit variance and different means. The three examples are: (1) 2 clusters, means of the sequences in cluster 1 and cluster 2 are $\{0.4,0.55,0.7,0.85,1,1.15,1.3,1.45,1.6\}$ and $\{1.85,2,2.15\}$, (2) 2 clusters, means of the sequences in cluster 1 and cluster 2 are $\{0.7,0.85,1,1.15,1.3\}$ and $\{1.7,1.85,2,2.15,2.3\}$, and (3) 5 clusters, 5 sequences per cluster, mean of each sequence in cluster $k$ is k1k-1 for $k=1,2,3,4,5$.
For the above examples, $d_{I}$ is equal to the maximum gap between two neighbouring distributions within a cluster, i.e, for each distribution we need only the left and right neighbours to be closer than the distributions in the other clusters. Therefore, we can have big clusters that are separated by gaps that are small (as in Example 1 and 2). In fact, for these examples, clustering is achieved by identifying the largest gap between means. Table II shows the MMD and KSD distances $d_{L}$, $d_{H}$, and $d_{I}$ for each of the three examples. The distances are estimated using 10000 samples from each distribution. [^9] Analytical evaluation of MMD is not straighforward and has been studied in [30] The MMD between distributions $p$ and $q$ can be calculated in terms of expected kernels for samples from $(p,p)$, $(q,q)$ and $(p,q)$. In Table 1 of [30], the analytical form of the expected kernel is provided some cases of $p$, $q$ and kernel function choices including the special case where $p$ and $q$ are Gaussian and the kernel is also Gaussian. We have checked that the values we report in Table II of our paper match with the distance evaluated using the expressions in [30]. Clearly, we have $d_{I}<d_{H}$ in all three examples, while $d_{L}<d_{H}$ is true only in example 3.
@@
Next, we consider two more examples with non-Gaussian data sequences. There are 2 clusters with 3 data sequences each. Here, each data sequence is generated in an i.i.d. manner according to a Gaussian Mixture Model (GMM). In particular, each sequence is a mixture of 2 Gaussian distributions: $N(m_{1},1)$ with probability $0.7$, and $N(m_{2},1)$ with probability 0.3. For Example 4, the $(m_{1},m_{2})$ values for the sequences in the two clusters are as follows: {(0.5,0),(0,0.5),(0.5,1)(-0.5,0),(0,0.5),(0.5,1)}, and {(1.2,1.7),(1.7,2.2),(2.2,2.7)(1.2,1.7),(1.7,2.2),(2.2,2.7)}. For Example 5, the $(m_{1},m_{2})$ values for the sequences in the two clusters are as follows: {(0.5,0),(0,0.5),(0.5,1)(-0.5,0),(0,0.5),(0.5,1)}, and {(1.35,1.85),(1.85,2.35),(2.35,2.85)(1.35,1.85),(1.85,2.35),(2.35,2.85)}. It can be seen that the clusters are closer in Example 4 compared to Example 5. The distances $d_{L}$, $d_{I}$, and $d_{H}$ for these two examples are given in Table III. In both cases, $d_{L}>d_{H}$, but $d_{I}<d_{H}$. In particular, $d_{I}$ is only slightly lesser than $d_{H}$ in Example 4.
Finally, we consider two real-world data set examples, one using Modified National Institute of Standards and Technology (MNIST) data [31], and the other using MovieLens data [32, 33]. The MNIST data set contains about 6000 images each of handwritten numerals from 0 to 9 of size $28\times 28$. Using this data set, we create five data sequences of 1000 samples for each numeral from 0 to 9. Thus, we have 50 data sequences forming 10 clusters. Here, each sample in a data sequence has $28\times 28=784$ dimensions For this multi-dimensional sample case, we use the kernel function:
[…]
We also created another lower dimensional data set that reduces the 784 dimensions to 10 dimensions using principal component analysis. The MovieLens data set consists of movie ratings from different viewers of movies from different genres. We take data of five genres and classify then into 3 clusters, i.e., 5 sequences and 3 clusters. The average ratings of the 5 genres: Horror, Adventure, Thriller, Romance and Crime are 3.2034, 3.4010, 3.4318, 3.4666 and 3.6014 respectively, corresponding to 3 clusters {Crime}, {Adventure, Thriller, Romance} and {Horror}.
@@ V-B FSS performance
Fig. 5 shows $\text{ln}(P_{e})$ versus the sample size $n$ for Example 1. The performance of SLINK is shown for both MMD (with kernel $k(x,y)=e^{-(x-y)^{2}/2}$) and KSD as the distance measures. It can be observed that SLINK is exponentially consistent, i.e., the $\text{ln}(P_{e})$ vs. $n$ plot is linearly decreasing, as expected. For this example, KMed is unable to find the true clusters in the FSS setting for any $n$. Although only the KSD case is shown for KMed, KMed with MMD also does not work. This is mainly due to the fact that one cluster is big while the other is small, and some distributions in the big cluster are closer to the medoid of the small cluster. However, SLINK can work even in this setting as $d_{I}<d_{H}$ is satisfied as long as the left and right neighbours for each distribution (in terms of the means of the Gaussian distributions) within the same cluster are closer than the distributions from the other clusters. In Table IV, we compare the observed slope in the simulation in Fig. 5 with the slope of the upper bound in Remark 1. We observe that the simulated performance slope is better than the upper bound by a factor about 25 in this case. The upper bound was mainly derived to show exponential consistency, and provides only a loose bound on the actual slope.
Figs. 6 and 7 show $\text{ln}(P_{e})$ versus the sample size $n$ for Examples 2 and 3. Both SLINK and KMed are exponentially consistent in the FSS setting for these examples. Even though $d_{L}>d_{H}$, KMed works for Example 2 since the clusters are of equal size and, therefore, all the distributions in each cluster are closer to the medoid of the correct cluster. For these two examples, KMed is able to perform better than SLINK, i.e., it achieves the same error probability at lower $n$. In particular, for example 2, KMed with MMD estimates can achieve $\text{ln}(P_{e})=-1.5$ with 100 samples, while SLINK with MMD estimates requires around 200 samples. However, SLINK can work for all problems with $d_{I}<d_{H}$, while KMed cannot. Fig. 7 also shows the performance of CLINK using both KSD and MMD estimates. It can be observed that for this example, CLINK is also exponentially consistent. It performs better than SLINK for lower $n$. However, CLINK does not work for Examples 1 and 2 since $d_{L}>d_{H}$ in these two examples. The comparison of the estimated slope from the simulation with the upper bound in Remark 1 is shown for these examples also in Table IV.
TABLE IV: Comparison of slope: Simulation vs. Analysis
Magnitude of slope of $\ln{P_{e}}$ vs. $n$
$b_{f}=\frac{(d_{H}-d_{I})^{2}}{64\mathbb{G}}$ in bound Estimated from Figs. 5-7
Example 1 $0.3773\times 10^{-4}$ $9.572\times 10^{-4}$
Example 2 $0.1694\times 10^{-3}$ $5.4885\times 10^{-3}$
Example 3 $0.2664\times 10^{-2}$ $9.084\times 10^{-2}$
@@ V-C Sequential clustering performance
Fig. 10 shows $\text{ln}(P_{e})$ versus $E[N]$ for SLINK-SEQ and SLINK FSS using MMD for the GMM Examples 4 and 5. Again, as in the previous examples, we observe that the proposed sequential clustering algorithm requires fewer expected number of samples than the FSS test for the same $P_{e}$. Finally, in Figs. 11 and 12, we show the performance of SLINK FSS and SLINK-SEQ for the MNIST and MovieLens data sets. Again, the algorithms are exponentially consistent. For the MNIST data, the proposed SLINK-SEQ algorithm is able to successfully cluster the data sequences corresponding to the same number. We also observe that the performance with the 784-dimensional data is better than the performance with the 10-dimensional data obtained using principal component analysis. For MovieLens data, we again observe exponential consistency and better performance of SLINK-SEQ compared to SLINK FSS.
@@
[…]
Figure 12: Performance improvement using SLINK-SEQ compared to SLINK FSS for MovieLens dataset, using MMD as distance
@@
In Figs. 13 and 14, we show the performance of SLINK-SEQ for different threshold choices for Examples 1, 2 and 3, respectively. We use the threshold $C/n^{\alpha}$ with $\alpha=1/3,1/2,1$. We observe that for all cases, $\alpha=1/2$ results in better performance.
@@ Appendix A Concentration of MMD estimates
$P\left(\mathbb{MMD}\left(\mathrm{x}_{1},\mathrm{x}_{2}\right)\leq d_{0}\right)\leq 2\exp\left(\frac{-n(d_{H}-d_{0})^{2}}{16\mathbb{G}}\right)$ (20)
@@
$P\left(\mathbb{MMD}\left(\mathrm{x}_{1},\mathrm{x}_{2}\right)\geq d_{0}\right)\leq 2\exp\left(\frac{-n(d_{0}-d_{I})^{2}}{16\mathbb{G}}\right)$ (22)